OpenStudy (anonymous):

a plane has take off speed of 88.3 m/s and requires 1365 m to reach that speed. determine the acceleration of the plane and the taken to reach the take off speed.

OpenStudy (anonymous):

You can get a start on this problem by using four symbols: t = time from start (at rest) to reaching takeoff speed a = acceleration of plane (assumed to be constant, which may be an unrealistic assumption for an actual plane, but is probably what the problem asker expects) v = takeoff speed of the plane x = distance from start point to where plane reaches takeoff speed Some equations you know from your course: v = a*t speed reached after t seconds at acceleration a meters/sec^2 x = (v/2)*t = (1/2)*a*t^2 distance reached after t seconds, at an average velocity of v/2 for t seconds, or at a constant acceleration a for t seconds (you can use either equation). And some numerical facts given to you in the problem: v = 88.3 meters/sec x = 1365 meters So now you can plug ithe v and x values nto the above equations, and get two equations involving unknowns a and t. Then figure out the values of a and t. Best wishes!